OpenStudy (christos):

4^2 - 4sin^2(x) = 4cos^2 true or false?

OpenStudy (anonymous):

sin^2+cos^2=1 if i remember correctly

OpenStudy (christos):

yep

OpenStudy (anonymous):

so all you have to do is test by moving the equations over.

OpenStudy (aakashsudhakar):

4^2 - 4sin^2(x) = 4cos^2(x) 4(2 - sin^2(x)) = 4(cos^2(x)) 2 - sin^2(x) = cos^2(x) sin^2(x) + cos^2(x) = 2 ^^ This is not true, therefore the equation you have is false.

OpenStudy (christos):

4 - 4sin^2(x) = 4cos^2 How about this ? True or false ?

OpenStudy (aakashsudhakar):

^^ That is true, because dividing the 4 from everything gets you... \[1 - sin^2(x) = cos^2(x)\] Which simplifies to... \[sin^2(x) + cos^2(x) = 1\] ^^ This is correct, therefore the equation you have is true.